What is the distance between E(-7,6), F(2,0)
Answer by stanbon(57940)

What is the distance between E(-7,6), F(2,0)
---
d = sqrt[(6-0)^2 + (-7-2)^2] = sqrt[36+81] = sqrt(117) = 10.82
===========================
Cheers,
Stan H.
==========